OpenStudy (anonymous):

how to solve for x ? 4^x - 2^x - 12 = 0

OpenStudy (anonymous):

\[4^{x} - 2^{x} - 12 = 0\]

hartnn (hartnn):

hint : take 2^x as y then 4^x = y^2 and you will get a quadratic in 'y' which you can solve easily :)

OpenStudy (anonymous):

so i will be getting y=4 and y=-3 . how do I find x when y= -3 ?

OpenStudy (anonymous):

@hartnn

OpenStudy (kira_yamato):

It would be rejected, unless you wanna use De Moivre's Theorem to get some complex number.

hartnn (hartnn):

lets stick with real domain and so y=4 2^x =4 can you find x from here ?

OpenStudy (anonymous):

x = 2 ? @hartnn

hartnn (hartnn):

thats it!

OpenStudy (anonymous):

so when y = -3 , x is invalid ?

OpenStudy (kira_yamato):

If and only if you stick to the real domain that is...

hartnn (hartnn):

yes, as 2^x cannot be negative for a real value of x its always >0